Another abbreviated status.  Andy should now be in Lisbon.  UPorto class tour today.  I'm still the "boss".

Quick look through cockpit shows no new problems:

  • TRH.20m.rsw06 still needs to be replaced
  • needs to be reprogrammed. Andy brought sensor to swap today.
  • sonic+li7500.v06 brought themselves back to life yesterday!
  • 4 ARL Li7500s should arrive in the next day or two
  • 3 tnw masts should be built today.  Now prepped in ops center
  • 3 tne masts should be started tomorrow, built by end of the week.  Prepping now... tse05 done.


  1. Steve Oncley AUTHOR

    About 1600, replaced with 940185, that Andy had brought.  Now, Rsw is sorted into in and out correctly.  It will be a pain to pick apart the data from this site up till now.

    I didn't have supplies with me, but these radiometers need to be cleaned, since the nearby surfaces have some kind of sticky coating.

  2. Steve Oncley AUTHOR

    For post-processing: I note that both and Rsw.out had ID 0x825a before this swap.  Thus, there will be no way to sort out the data using the standard ID numbers.  It probably will be necessary to write code to look at the values, decide which is which, create appropriate averages, and write to 5-min NetCDF.  This should be straightforward, since both data samples will have the same time tag and should appear in the same order:

    2017 05 01 16:00:04.8492       0 35,0x825a       4      590.2

    2017 05 01 16:00:04.8492       0 35,0x825a       4       61.5 

    2017 05 01 16:00:09.8487       0 35,0x825a       4      589.9 

    2017 05 01 16:00:09.8487       0 35,0x825a       4       61.4 

    The new data start at 15:50.